### v2

@@ 2.1 Data
The frequencies above differ from frequencies used in traditional corpus studies in the sense that they are calculated on the preferred orders of languages. Consider the frequency of nAND in languages, that is 182 according to (Dryer, 2018); see Table 2.1. This frequency is the number of languages that prefer the order nAND (and not any of the other 23 possibilities) for a noun phrase consisting of a demonstrative, a numeral, an adjective and a noun. It is not the frequency of that order in a corpus, although one would expect nAND to be the most frequent in a corpus in each of the 182 languages.
@@ 2.2 The models
The term power law is used both for continuous and discrete random variables (Debowski, 2020; Stumpf & Porter, 2012; Newman, 2005; Conrad & Mitzenmacher, 2004; Naranan & Balasubrahmanyan, 1998). Similarly, the term exponential distribution is used both for continuous and discrete random variables (Broido and Clauset, 2019; Clauset et al., 2009; Newman, 2005; Ferrer-i-Cancho, 2004). In these contexts, the term power law is used to refer to a distribution that approximates a power-law function. This is why often the term power-law-like distribution is used. Similarly, the term exponential is used to refer to a distribution that approximates an exponential function. Analogously, one can also use the term exponential-like distribution. However, here our random variable, i.e. rank, is discrete. Martin et al. (2020) stated that the distribution is a power law but did not specify the actual form of the distribution. Our first task is to translate informal terminology into specific discrete distributions (Wimmer & Altmann, 1999; Johnson et al., 2005).
@@ 3.1 Visual diagnostic
Figure 2: $f(r)$, the frequency of a word order of rank $r$ in Dryer (2006) in three distinct scales: normal (top), linear-log (middle) and log-log (bottom). The solid line is the real curve and the discontinuous lines are the expected value of $f(r)$, that is $\mathbb{E}[f(r)]=F_{0}p(r)$, where $p(r)$ is defined by the best fit of a model (Table 3.2). Left. Real curves versus the best fit of the Zeta 1 model (dotted line) and that of Zeta 2 model (dashed line). Right. Real curves versus the best fit of Geometric 1 (dotted line) and that of Geometric 2 (dashed line).
@@ 4.1.2 The correct exponential distribution
We have found that the best model depends on the score. According to $AIC_{c}$, Geometric 2 has more evidence than Geometric 1 (except for the 2006 dataset). According to $BIC$, Geometric 1 has more evidence (except for adjusted number of languages in the 2018 dataset). However, a stronger conclusion can be reached by looking at the ability of the models to generalize in perspective. Cinque (2005) reported only 14 attested orders. Although he did not report the frequencies of each attested on a sample, we can be totally certain that the best fit of Geometric 2 by maximum likelihood would conclude that only $R=14$ orders have non-zero probability (justification in Appendix C), which we know it would be wrong because 17 and 18 orders were found later on (Dryer, 2006, 2018). We also know that the best fit of Geometric 2 to the 2006 dataset, namely $R=17$ is misleading because $R=18$ for the 2018 dataset (Table 3.2). Thus, Geometric 2 fails to generalize two times.
@@
When integrating likelihood into an information theoretic criterion and considering again the 2006 dataset, $AIC_{c}$ weights conclude that Geometric 1 is as likely as Geometric 2 (recall $w_{2}(AIC_{c})/w_{2}(AIC_{c})\approx 1$) but $BIC$ weights are able to realize that Geometric 1 is more likely to be the best (recall $w_{1}(BIC_{c})/w_{2}(BIC)\approx 6$). Crucially, $AIC_{c}$ does not foresee that the best model in the 2006 dataset, Geometric 2, produces a zero likelihood when applied to the 2018 dataset (notice that the application of the best fit of Geometric 2 for the 2006 dataset to the 2018 dataset yields $p(18)=0$, which produces $L=0$ when applied to equation 9). In contrast, the best model for the 2006 dataset according to BIC yields a non-zero likelihood when applied to the 2018 dataset. In sum, BIC catches early the model that generalizes while $AIC_{c}$ is unable to realize that one of the models overfits the 2006 dataset. Therefore, $AIC_{c}$ lacks generalizability.
@@
The failure of $AIC_{c}$ on the 2006 dataset is not surprising given the theoretical properties of $AIC_{c}$ versus $BIC$ (Wagenmakers & Farrell, 2004). $AIC$ is not consistent in the sense that, as the number of observations ($F_{0}$ in our notation) grows very large, the probability that $AIC_{c}$ recovers a true low-dimensional model does not converge to $1$ (Bozdogan, 1987, p. 357). $BIC$ is consistent as the number of observations tends to infinity. In our application, that means that if we supplied to $AIC_{c}$ a dataset comprising all languages on Earth or even all languages that ever existed in our planet, it would not be warranted that $AIC_{c}$ would recover the right geometric models while $BIC$ would be much closer to finding the right geometric model. $BIC$ is more useful in selecting a correct model (in this case, discarding an incorrect model) while the AIC is more appropriate in finding the best model for predicting future observations. (Chakrabarti and Ghosh, 2011). In our application, the main goal is to find the right geometric model, not a geometric model that predicts the actual shape of the distribution when future observations are incorporated.
@@
Above, we have provided empirical evidence of the failure of $AIC_{c}$ to find a correct model in our application. However, we wish to highlight the theoretical power of $BIC$ for our research problem and our datasets. Once we have discarded the power-law models, the model selection reduces to choosing between two geometric models. $BIC$ assumes equal prior probability on each model (Burnham and Anderson, 2002). That implies assigning initial equal chance to early truncation (Geometric 2) and to late truncation (Geometric 1). Indeed, such a contest between two nested models, Geometric 1 and Geometric 2, can be seen as a dimensionality guessing problem (Geometric 2 adds on dimension with respect to Geometric 1) and $BIC$ is a consistent estimator of $K$, the dimensions of the “true model” (Burnham and Anderson, 2002, p. 284). $BIC$ implicitly assumes that “truth is of fairly low dimension (e.g., K=15K=1-5) and that the data-generating (true) model is fixed as sample size increases” (Burnham and Anderson, 2002, 286). Various sources of evidence suggest that the true model’s dimensionality is bounded by a small number in our application. Dryer (2018) accurately modelled the frequency rank of the orders by means of 5 binary descriptive principles, suggesting that the true number of dimensions of the distribution is bounded above by $K=5$. Indeed, just $\lfloor\log_{2}N\rfloor+1=5$ binary parameters suffice to sort $N$ orders so as to match the desired frequency rank. Previously, Cysouw (2010) had proposed exponential models for the frequency of orders with 3, 4, or 6 binary parameters. Therefore, we apply BIC under favourable conditions, where the dimensions of truth are bounded. In our application, $BIC$ always provides more evidence for Geometric 1 except for adjusted number of languages in the 2018 dataset, that coincides with the smallest sample (Table 2.1). The consistency property of $BIC$ requires a large enough sample (Equation 8 and Burnham and Anderson (2002, p. 288)). Therefore, our findings so far and the conditions where we are applying $BIC$ suggest that Geometric 1 is a stronger model for the underlying distribution.
@@ 4.2 Are there hard constraints on word order?
Once the power law models are discarded, the question of the existence of hard constraints reduces to a dimensionality guessing problem, i.e. which of the two geometric models is the best. The evidence for Geometric 1 is a challenge for the existence of a hard constraints limiting word order variation in languages that would explain why not all 24 possible orders are attested (Cinque, 2005, 2013; Medeiros et al., 2016; Medeiros, 2018). If there were no hard constraints (only soft constraints), the best model should be Geometric 1; if strong constraints existed, Geometric 2 (with $R<24$) should be the best model. The strength of Geometric 1 is consistent with Cysouw’s hypothesis: all orders are a priori possible but some are not attested due to undersampling (Cysouw, 2010). The challenge of proponents of a hard constraint is two-fold. First, to make a robust proposal, one that does not fall any time that new orders are attested. For instance, the proposal that the 14 orders attested by Cinque (2005) result from universal grammar or some universal cognitive mechanism, still stands as soft constraint but not as hard constraint (Cinque, 2005, 2013; Medeiros et al., 2016; Medeiros, 2018) because so far, 18 orders have been attested (Dryer, 2018). Model selection by means of $AIC_{c}$ commits overfitting for believing that the currently number of attested orders is the true one. In particular, $AIC_{c}$ leads to parameters for the best model on the data from Dryer (2006) that do not generalize to the data from Dryer (2018). Second, to deal with parsimony. Postulating a hard constraint implies a loss of parsimony but is it rewarding enough in terms of proximity to truth? Information criteria address this question and give a compelling answer: when empirical generalizability and theoretical consistency are prioritized (this is the virtue of $BIC$ with respect to $AIC_{c}$), the absence of a hard constraint (Geometric 1) becomes more likely than its presence (Geometric 2), as shown in Table LABEL:tbl:model_selection_summary).
It is important to understand that answer to the question on the existence of hard constraints does not follow from choosing BIC because it penalizes for lack of parsimony more strongly than AIC (Wagenmakers & Farrell, 2004). If we did so, the conclusion that Geometric 1, and thus the absence of hard constraints, would be trivial. The relevant questions, that summarize the discussions above, are the following:
@@ 5 Conclusion
We have shown that a geometric distribution gives a better fit than a power law distribution to the distribution of preferred orders in the noun phase. This has implications for various components of the generative-linguistics program and other branches of theoretical linguistics the share the same abstractions, e.g., binary acceptability,  [^3] The point is assuming that there are impossible orders; whether acceptable orders vary in degree of acceptability is irrelevant for this point. or the same methods. On the one hand, our analysis shows that statistical evidence for a hard constraint on word order variation in the noun phrase is lacking. Not observing the 24 orders is not enough to claim for a hard constraint. History has demonstrated that the post-hoc arguments built on the current number of attested orders (Cinque, 2005, 2013; Medeiros et al., 2016; Medeiros, 2018) have collapsed as more orders have been attested. On the other hand, the finding of an exponential distribution challenges another component of the generative-linguistics program, i.e. the inevitability of power laws such as Zipf’s rank-frequency law, and consequently, the lack of interest in explaining their origin (Miller & Chomsky, 1963). Interestingly, linguistic laws are absent from the agenda in the continuum between generative linguistics and their opponents. When claiming that linguistic universals are a myth, opponents ignore their existence or neglect them (Evans & Levinson, 2009). An intriguing question is whether they do it for the same reason as generativists and formal linguists from other traditions. While researchers engage in complex debates about the potential advantage of models with a huge number of parameters (Futrell & Mahowald, 2025) as if reality were intrinsically of ultra-high dimension, the examination of elementary exponential distributions with a few parameters reveals that a hard constraint on word order lacks statistical support, as we have seen here, while it sheds light on the structure of short-term memory and the dynamics of incremental sentence processing (Petrini & Ferrer-i-Cancho, 2025). Reality may be simple, but researchers may fail to see it.

### v3

@@ 2.1 Data
The frequencies above differ from frequencies used in traditional corpus studies in the sense that they are calculated on the preferred orders of languages. Consider the frequency of nAND in languages, that is 182 according to Dryer (2018) (Table 2.1). This frequency is the number of languages that prefer the order nAND (and not any of the other 23 possibilities) for a noun phrase consisting of a demonstrative, a numeral, an adjective and a noun. It is not the frequency of that order in a corpus, although one would expect nAND to be the most frequent in a corpus in each of the 182 languages.
@@ 2.2 The models
The term power law is used both for continuous and discrete random variables (Debowski, 2020; Stumpf & Porter, 2012; Newman, 2005; Conrad & Mitzenmacher, 2004; Naranan & Balasubrahmanyan, 1998). Similarly, the term exponential distribution is used both for continuous and discrete random variables (Broido and Clauset, 2019; Clauset et al., 2009; Newman, 2005; Ferrer-i-Cancho, 2004). In these contexts, the term power law is used to refer to a distribution that approximates a power-law function. This is why the term power-law-like distribution is often used. Similarly, the term exponential is used to refer to a distribution that approximates an exponential function. Analogously, one can also use the term exponential-like distribution. However, here our random variable, i.e. rank, is discrete. Martin et al. (2020) stated that the distribution is a power law but did not specify the actual form of the distribution. Our first task is to translate informal terminology into specific discrete distributions (Wimmer & Altmann, 1999; Johnson et al., 2005).
@@ 3.1 Visual diagnostic
Figure 2: $f(r)$, the frequency of a word order of rank $r$ in Dryer (2006), in three distinct scales: normal (top), linear-log (middle) and log-log (bottom). The solid line is the real curve and the discontinuous lines are the expected value of $f(r)$, that is $\mathbb{E}[f(r)]=F_{0}p(r)$, where $p(r)$ is defined by the best fit of a model (Table 3.2). Left. Real curves versus the best fit of the Zeta 1 model (dotted line) and that of Zeta 2 model (dashed line). Right. Real curves versus the best fit of Geometric 1 (dotted line) and that of Geometric 2 (dashed line).
@@ 4.1.2 The correct exponential distribution
We have found that the best model depends on the score. According to $AIC_{c}$, Geometric 2 has more evidence than Geometric 1 (except for the 2006 dataset). According to $BIC$, Geometric 1 has more evidence (except for adjusted number of languages in the 2018 dataset). However, a stronger conclusion can be reached by looking at the ability of the models to generalize in perspective. Cinque (2005) reported only 14 attested orders. Although he did not report the frequencies of each attested on a sample, we can be totally certain that the best fit of Geometric 2 by maximum likelihood would conclude that only $R=14$ orders have non-zero probability (justification in Appendix C), which we know it would be wrong because 17 and 18 orders were found later on (Dryer, 2006, 2018). We also know that the best fit of Geometric 2 to the 2006 dataset, namely $R=17$ is misleading because $R=18$ for the 2018 dataset (Table 3.2). Thus, Geometric 2 fails to generalize twice.
@@
When integrating likelihood into an information theoretic criterion and considering the 2006 dataset again, $AIC_{c}$ weights conclude that Geometric 1 is as likely as Geometric 2 (recall $w_{2}(AIC_{c})/w_{2}(AIC_{c})\approx 1$) but $BIC$ weights are able to realize that Geometric 1 is more likely to be the best (recall $w_{1}(BIC_{c})/w_{2}(BIC)\approx 6$). Crucially, $AIC_{c}$ does not foresee that the best model in the 2006 dataset, Geometric 2, produces a zero likelihood when applied to the 2018 dataset (notice that the application of the best fit of Geometric 2 for the 2006 dataset to the 2018 dataset yields $p(18)=0$, which produces $L=0$ when applied to equation 9). In contrast, the best model for the 2006 dataset according to BIC yields a non-zero likelihood when applied to the 2018 dataset. In sum, BIC catches early the model that generalizes while $AIC_{c}$ is unable to realize that one of the models overfits the 2006 dataset. Therefore, $AIC_{c}$ lacks generalizability.
@@
The failure of $AIC_{c}$ on the 2006 dataset is not surprising given the theoretical properties of $AIC_{c}$ versus $BIC$ (Wagenmakers & Farrell, 2004). $AIC$ is not consistent in the sense that, as the number of observations ($F_{0}$ in our notation) grows very large, the probability that $AIC_{c}$ recovers a true low-dimensional model does not converge to $1$ (Bozdogan, 1987, p. 357). $BIC$ is consistent as the number of observations tends to infinity. In our application, that means that if we supplied to $AIC_{c}$ a dataset comprising all languages on Earth or even all languages that ever existed in our planet, it would not be warranted that $AIC_{c}$ would recover the right geometric models while $BIC$ would be much closer to finding the right geometric model. $BIC$ is more useful in selecting a correct model (in this case, discarding an incorrect model) while the AIC is more appropriate in finding the best model for predicting future observations (Chakrabarti and Ghosh, 2011). In our application, the main goal is to find the right geometric model, not a geometric model that predicts the actual shape of the distribution when future observations are incorporated.
@@
Above, we have provided empirical evidence of the failure of $AIC_{c}$ to find a correct model in our application. However, we wish to highlight the theoretical power of $BIC$ for our research problem and our datasets. Once we have discarded the power-law models, the model selection reduces to choosing between two geometric models. $BIC$ assumes equal prior probability on each model (Burnham and Anderson, 2002). That implies assigning initial equal chance to early truncation (Geometric 2) and to late truncation (Geometric 1). Indeed, such a contest between two nested models, Geometric 1 and Geometric 2, can be seen as a dimensionality guessing problem (Geometric 2 adds on dimension with respect to Geometric 1) and $BIC$ is a consistent estimator of $K$, the dimensions of the “true model” (Burnham and Anderson, 2002, p. 284). $BIC$ implicitly assumes that “truth is of fairly low dimension (e.g., K=15K=1-5) and that the data-generating (true) model is fixed as sample size increases” (Burnham and Anderson, 2002, 286). Various sources of evidence suggest that the true model’s dimensionality is bounded by a small number in our application. Dryer (2018) accurately modelled the frequency rank of the orders by means of 5 binary descriptive principles, suggesting that the true number of dimensions of the distribution is bounded above by $K=5$. Indeed, just $\lfloor\log_{2}N\rfloor+1=5$ binary parameters suffice to sort $N$ orders so as to match the desired frequency rank. Previously, Cysouw (2010) had proposed exponential models for the frequency of orders with 3, 4, or 6 binary parameters. Therefore, we apply BIC under favourable conditions, where the dimensions of truth are bounded. In our application, $BIC$ always provides more evidence for Geometric 1 except for adjusted number of languages in the 2018 dataset, which coincides with the smallest sample (Table 2.1). The consistency property of $BIC$ requires a large enough sample (Equation 8 and Burnham and Anderson (2002, p. 288)). Therefore, our findings so far and the conditions where we are applying $BIC$ suggest that Geometric 1 is a stronger model for the underlying distribution.
@@ 4.2 Are there hard constraints on word order?
Once the power law models are discarded, the question of the existence of hard constraints reduces to a dimensionality guessing problem, i.e. which of the two geometric models is the best. The evidence for Geometric 1 is a challenge for the existence of hard constraints limiting word order variation in languages that would explain why not all 24 possible orders are attested (Cinque, 2005, 2013; Medeiros et al., 2016; Medeiros, 2018). If there were no hard constraints (only soft constraints), the best model should be Geometric 1; if strong constraints existed, Geometric 2 (with $R<24$) should be the best model. The strength of Geometric 1 is consistent with Cysouw’s hypothesis: all orders are a priori possible but some are not attested due to undersampling (Cysouw, 2010). The challenge of proponents of a hard constraint is two-fold. First, to make a robust proposal, one that does not fall any time that new orders are attested. For instance, the proposal that the 14 orders attested by Cinque (2005) result from universal grammar or some universal cognitive mechanism, still stands as soft constraint but not as hard constraint (Cinque, 2005, 2013; Medeiros et al., 2016; Medeiros, 2018) because so far, 18 orders have been attested (Dryer, 2018). Model selection by means of $AIC_{c}$ commits overfitting for believing that the current number of attested orders is the true one. In particular, $AIC_{c}$ leads to parameters for the best model on the data from Dryer (2006) that do not generalize to the data from Dryer (2018). Second, to deal with parsimony. Postulating a hard constraint implies a loss of parsimony but is it rewarding enough in terms of proximity to truth? Information criteria address this question and give a compelling answer: when empirical generalizability and theoretical consistency are prioritized (this is the virtue of $BIC$ with respect to $AIC_{c}$), the absence of a hard constraint (Geometric 1) becomes more likely than its presence (Geometric 2), as shown in Table LABEL:tbl:model_selection_summary).
It is important to understand that the answer to the question on the existence of hard constraints does not follow from choosing BIC because it penalizes for lack of parsimony more strongly than AIC (Wagenmakers & Farrell, 2004). If we did so, the conclusion that Geometric 1, and thus the absence of hard constraints, would be trivial. The relevant questions, that summarize the discussions above, are the following:
@@ 5 Conclusion
We have shown that a geometric distribution gives a better fit than a power law distribution to the distribution of preferred orders in the noun phase. This has implications for various components of the generative-linguistics program and other branches of theoretical linguistics that share the same abstractions, e.g., binary acceptability,  [^3] The point is assuming that there are impossible orders; whether acceptable orders vary in degree of acceptability is irrelevant for this point. or the same methods. On the one hand, our analysis shows that statistical evidence for a hard constraint on word order variation in the noun phrase is lacking. Not observing the 24 orders is not enough to claim a hard constraint. History has demonstrated that the post-hoc arguments built on the current number of attested orders (Cinque, 2005, 2013; Medeiros et al., 2016; Medeiros, 2018) have collapsed as more orders have been attested. On the other hand, the finding of an exponential distribution challenges another component of the generative-linguistics program, i.e. the inevitability of power laws such as Zipf’s rank-frequency law, and consequently, the lack of interest in explaining their origin (Miller & Chomsky, 1963). Interestingly, linguistic laws are absent from the agenda in the continuum between generative linguistics and its opponents. When claiming that linguistic universals are a myth, opponents ignore their existence or neglect them (Evans & Levinson, 2009). An intriguing question is whether they do it for the same reason as generativists and formal linguists from other traditions. While researchers engage in complex debates about the potential advantage of models with a huge number of parameters (Futrell & Mahowald, 2025) as if reality were intrinsically of ultra-high dimension, the examination of elementary exponential distributions with a few parameters reveals that a hard constraint on word order lacks statistical support, as we have seen here, while it sheds light on the structure of short-term memory and the dynamics of incremental sentence processing (Petrini & Ferrer-i-Cancho, 2025). Reality may be simple, but researchers may fail to see it.